If a solution containing 57.620 g of mercury(II) perchlorate is allowed to react completely with a solution containing 15.488 g of sodium dichromate, how many grams of solid precipitate will be formed?
How many grams of the reactant in excess will remain after the reaction?
1
Expert's answer
2019-04-05T07:23:53-0400

 X mol       0.059 mol     Y mol

Hg(ClO4)2 + Na2Cr2O7 = HgCr2O7↓ + 2NaClO4

 1 mol           1 mol          1 mol


n(Hg(ClO4)2) = m(Hg(ClO4)2)/M(Hg(ClO4)2) = 57.620 g / 400 g/mol = 0.144 mol – excess.

n(Na2Cr2O7) = m(Na2Cr2O7)/M(Na2Cr2O7) = 15.488 g / 262 g/mol = 0.059 mol – deficiency.


n(HgCr2O7) = Y = 0.059 mol.

m(HgCr2O7) = M(HgCr2O7)×n(HgCr2O7) = 417 g/mol × 0.059 mol = 24.603 g.

nexc(Hg(ClO4)2) = n(Hg(ClO4)2) – X = 0.144 mol – 0.059 mol = 0.085 mol.

mexc(Hg(ClO4)2) = M(Hg(ClO4)2)×nexc(Hg(ClO4)2) = 400 g/mol × 0.085 mol = 34 g.